# Autograding Mathematical Induction Proofs with Natural Language Processing

### v1

@@ 4 Model Performance Results
In this section, we will provide an overview on the performance of the grading models, the accuracy of the human graders, and an analysis of how much training data we need to get satisfactory performance.
@@ 4.1 Grading Model Performance
For each model and each training epoch, we calculate the performance using the validation set, comparing the model predictions to ground truth labels that previous researchers created. We compute the confusion matrix to obtain the number of true positive, true negative, false positive, false negative cases, and to calculate the accuracies and F1 scores. Among the models using different training epochs, we select the one with the highest accuracy. Model results for each problem and each pretrained model are summarized by averaging the accuracies from the testing set for all rubric points. The results are shown in Table 2.
Problem Data Size Grading Method Accuracy
P1 1623 MathBERT & linear 83.6%
GPT-3 & linear 87.9%
Llemma7b & linear 90.7%
Llemma34b & linear 90.2%
Human Graders 84.1%
P2 1288 MathBERT&linear 83.7%
GPT-3 & linear 88.3%
Llemma7b & linear 90.9%
Llemma34b & linear 90.7%
Human Graders 91.3%
P3 342 MathBERT&linear 85.7%
GPT-3 & linear 89.2%
Llemma7b & linear 85.7%
Llemma34b & linear 86.6%
Human Graders 85.1%
P4 333 MathBERT&linear 86.3%
GPT-3 & linear 85.4%
Llemma7b & linear 89.9%
Llemma34b & linear 90.2%
Human Graders 85.8%
@@
All four pretrained models achieve over $80\%$ average accuracy. MathBERT is the lowest performing one with an average of $84.1\%$ accuracy across all problems, as it is the model with the fewest parameters, thus capturing fewer details accurately. The best grading models are Llemma-based, both Llemma34b and Llemma7b, with $90.0\%$ and $90.2\%$ average accuracy, even beating GPT-3-based models. This is likely because Llemma is more specifically trained to handle mathematical language. Moreover, the large number of parameters gives Llemma the ability to capture almost the same amount of information as GPT-3. The accuracies and F1 scores of Llemma34b graders on a by-rubric-point basis appears in Table 3.
@@ A F1 A F1 A F1 A F1 A F1 A F1 A F1
Based on these results, there seems to be no pattern of which rubric points are more easily gradable than others. Currently, Llemma and Llemma-based grading models perform well enough for use by students as practice for writing mathematical proofs. In future work, we want to investigate how to provide more accurate and detailed feedback to students, as the correctness of the models is not guaranteed.
@@ 4.3 How large of a dataset do we need?
For all four models, the performance is still increasing even with 1100 training data points, so there is space for improvement by collecting more data. Model performance might also be better by making the grading models more complex, such as adding extra layers. However, most of the improvement seems to be complete before the 400 training size threshold. For Llemma34b, the performance at 400 training data already achieves 98.8% of the full data size accuracy. Based on these results, we conclude that training a model with 400 proofs is sufficient for adequate results.
@@ 5 User Study Methods
Using the trained grading pipelines, we implemented an autograder for proof by induction problems using the online assessment system PrairieLearn [35]. Due to computational resource restrictions and the fact that GPT-3 performed nearly as well as Llemma on the 3 proofs we used for the user study, we chose to use GPT-3 as the embedding model for our user study. For each submitted student proof on a problem, the grading pipeline reports seven correct/incorrect results using the rubrics R1-R7 introduced in 3.1. The final score for the proof is the sum of the 7 rubric point results, ranged between 0 and 7, and the rescaled to 0-100. The grader then provides feedback on the rubric point results based on various strategies we defined. Figure 4 shows an example page from PrairieLearn where students can submit their proofs and receive automatic feedback.
@@ 5.1 Study Setup
The learning activity consisted of three parts. In the first part, students were expected to read a 6-page textbook section about induction, including information on the theoretical ground for induction, the steps of an induction proof, and an example problem with solution. The second part asked students to complete 3 induction problems (P1-P3 as defined in 3.1). We only consider 3 problems out of the 4 used in the model training process for two reasons: first, we wanted to limit the length of the learning activity to 1 hour, which would not be reasonable if we expected students to read the text and complete all 4 problems; second, we removed P4 due to its low accuracy compared to P1-P3, as this could potentially hinder the students’ learning experiences. The third part of the learning activity was a feedback survey. The questions are listed below:
@@
The first subsection of the survey includes questions S01-S03, which asks the students about their prior experience with human graders in the course. The second subsection includes questions S04-S07, focusing on students’ perceptions of using AI for autograding. The last questions ask students to share their experiences with the AI autograder during the learning actvity. The survey questions are presented on a five-point Likert scale using the following options and corresponding numeric values: “Strongly disagree” (-2), “Disagree” (1), “Neutral” (0), “Agree” (1), “Strongly agree” (2). Negatively worded questions (e.g., “I would prefer to wait for a week to get human grading feedback for the proofs I write for my homework than use the autograder for instant feedback”) are reverse coded in the statistical analysis.
To help us get a more full picture of student perceptions about the graders, we also asked the following open-ended survey questions:
@@ 5.2 Experimental Conditions
Both the First and Random groups had access to the AI autograders, forming our treatment groups. Students in the First group received feedback on the first rubric point out of the seven that the autograder identified them having done incorrectly. The order is defined by the sequence of rubrics R1-R7, which corresponds to the logical flow of a proof. For an induction proof, it is common to state and prove the base cases first, then state the goal for the induction steps before actually completing the steps. The strategy of reporting the first incorrect rubric point is based on some students’ preference to work incrementally, fixing the earlier parts of their work before moving on to the next part. When we designed the experiment, we expected this strategy to be the most helpful to students.
Students in the Random group received feedback for one of rubric points, randomly selected from all those that the autograder assessed as incorrect. This group is included in our study to test if different strategies for reporting have different effects.
@@ 6 User Study Results
6.1 Proof Problem Performance
@@ 6.2 Survey Results
Figure 8: Responses from survey questions given to only First and Random. The percentages shown on each distribution represent the percentage of negative responses (disagree & strongly disagree), the percentage of netural responses, and the percentage of positive responses (agree & strongly agree).
The survey results are shown in Figure 7 and Figure 8. The three subsections of the surveys address different aspect of student perceptions: S01-S03 address student perceptions of human graders, S04-S07 address students willingness to have AI graders integrated into various parts of their course, and S08-S10 address student perceptions of the AI graders. Wee calculated a Cronbach’s $\alpha$ for each of the three subsections. The calculated Cronbach’s $\alpha$ are 0.81, 0.72, 0.82 respectively, indicating good internal reliability. The Self-eval group were not asked questions S08-S14 as these students did not interact with the autograder during the learning activity.
@@ 6.2.1 RQ2: Are students satisfied with the autograder and feedback?
We use responses to survey questions S01-S03 and S08-S10, as shown in Table  7, to address RQ2. S01-S03 capture students’ perceptions of human graders, while S08-S10 assess their perceptions of the AI autograder. Specifically, S01 and S08 measure the perceived accuracy of the graders, S02 and S09 evaluate their helpfulness, and S03 and S10 measure overall satisfaction with the graders. On average, human graders are perceived to be more accurate and helpful than the AI autograder, with statistically significant differences ($p<0.05$). However, although the mean satisfaction rating is higher for human graders, the differences are not statistically significant for either the First or Random groups. We hypothesize that this is due to the known benefits of the AI autograder, such as timely feedback and the ability to submit an unlimited number of times.
@@ 6.2.2 RQ3: Does using the autograder make students more willing to use AI autograders?
We want to understand students’ perception and trust on using an autograder for their coursework during various settings. We want to know whether interactions with the autograder can lead to more positive perceptions. These are reflected through S04-S07 (see Table 8). S04 examines students’ preferences between receiving instant feedback from AI autograders and obtaining more reliable feedback from human graders. S05 assesses the level of trust students have in AI autograders in general. S06 and S07 explore potential applications of AI autograders.
@@ 7.2 Feedback Helpfulness
Despite our findings showing that students are able to make improvements using the feedback from the autograder, some students felt that the feedback was not detailed enough. The outputs from the grading models are always booleans representing correctness: they can tell if a student proof is incorrect, but not why the proof is incorrect. This form of feedback roughly aligns with knowledge of results defined by Shute [6], which was shown to be less effective than elaborated feedback in several studies [36]. The lack of useful feedback that can help students improve their proofs is also reported.
  • “Some of the feedback was basic. didn’t say anything specific. ”
Providing such automated and elaborated feedback should be possible with the help of generative models. Using few-shot training techniques such as Retrieval-Augmented Generation [37], we can fine-tune chatbots such as ChatGPT to provide suggestions or hints that can help students make improvements and are better aligned with the learning requirements for individual courses. This enhancement to our system can be treated in future work.
On the other hand, some students acknowledged that having the instant feedback was already very helpful in their proof-writing process.
@@ 7.3 Perceptions on AI grading
Proof grading is a subjective task, which can lead to low levels of trust from students. In our initial designing of the learning activities, we had hoped that students could have more trust or willingness in using AI autograders in the future as a result of interacting with our grader. However, results in 6.2 do not confirm our initial hypothesis. Students in First and Random are slightly more willing to use an AI autograder in various settings, but the difference is not statistically significant. Simply using the autograder does not sufficiently increase students’ willingness to use AI autograders in general. One possible action to enhance trust in AI autograding is to provide more feedback and explanation during the grading process, as Ha and Kim suggests that providing additional information about AI models can be effective in mitigating cognitive bias [38].
@@ 8 Conclusions and Future Work
In this study, we have developed a set of machine learning models to autograde freeform mathematical proofs. These models allow for fast training, and they achieve similar accuracy in autograding freeform mathematical proofs compared with most human graders recruited in the study. These models can be used to develop NLP graders and deployed on large-scale educational platforms in the future, enabling more students to receive the necessary support when learning to write mathematical proofs.
After creating and deploying autograded proof by induction problems using the grading models, we conducted a user study to find the impact of the autograder. Our results show that students who use the autograder for getting feedback and making improvements were able to score more than 10% higher on the proof problems than students who did not use the autograder. This is an encouraging sign for creating more autograded problems in the future. Despite being helped by the autograder, many students still preferred feedback from human graders, even if it was delayed.
In the future, it can be helpful to put effort in enabling more student trust in AI autograders. Efforts can be put on improving the grading models as well. As we only trained the grading models on mathematical induction problems, the abilities for similar models to grade other types of proof such as proof of bijection or geometry is still unknown. We will start to collect proof data for various types of mathematical proof problems, and check if our designed grading models can perform as well on these problems. The cost of training and running the models can also be lowered, as currently the best performing models require good GPUs to run, and the CPU-based models do not achieve performances that are as good.

### v2

@@ 4 Model Performance Results
In this section, we provide an overview of the grading models’ performance, the accuracy of the human graders, and an analysis of the amount of training data required to get satisfactory performance.
@@ 4.1 Grading Model Performance
For each model and each training epoch, we calculate the performance using the testing set, comparing the model predictions to ground truth labels that previous researchers created. We compute the confusion matrix to obtain the number of true positive, true negative, false positive, false negative cases, and then we calculate the accuracies and F1 scores. Among the models using different training epochs, we select the one with the highest accuracy. Rubric point accuracy metrics for each problem and each pretrained model are summarized by averaging the accuracies from the testing set for all rubric points. RMSE and Pearson’s $r$ are calculated by adding up the results for each rubric point, and normalizing the grades to between 0 and 100. The results are shown in Table 2.
Problem Data Size Grading Method Accuracy RMSE Pearson’s $r$
P1 1623 MathBERT & linear 83.6% 23.35 0.79
GPT-3 & linear 87.9% 17.06 0.90
Llemma7b & linear 90.7% 14.53 0.92
Llemma34b & linear 90.2% 15.33 0.92
Human Graders 84.1% 22.20 0.81
P2 1288 MathBERT&linear 83.7% 27.88 0.77
GPT-3 & linear 88.3% 22.32 0.86
Llemma7b & linear 90.9% 16.28 0.92
Llemma34b & linear 90.7% 17.87 0.91
Human Graders 91.3% 13.58 0.94
P3 342 MathBERT&linear 85.7% 20.51 0.85
GPT-3 & linear 89.2% 17.07 0.88
Llemma7b & linear 85.7% 21.98 0.81
Llemma34b & linear 86.6% 18.59 0.87
Human Graders 85.1% 23.55 0.78
P4 333 MathBERT&linear 86.3% 21.13 0.76
GPT-3 & linear 85.4% 17.86 0.83
Llemma7b & linear 89.9% 13.68 0.91
Llemma34b & linear 90.2% 13.98 0.91
Human Graders 85.8% 25.56 0.71
@@
All four pretrained models achieve over $80\%$ average accuracy, and achieve higher RMSE and Pearson’s $r$ metrics than most of the earlier ASAG work [33]. MathBERT is the lowest performing one with an average of $84.1\%$ accuracy across all problems, as it is the model with the fewest parameters, thus capturing fewer details accurately. The best grading models are Llemma-based, both Llemma34b and Llemma7b, with $90.0\%$ and $90.2\%$ average accuracy, even beating GPT-3-based models. This is likely because Llemma is more specifically trained to handle mathematical language. Moreover, the large number of parameters gives Llemma the ability to capture almost the same amount of information as GPT-3. The accuracies and F1 scores of Llemma34b graders on a by-rubric-point basis appears in Table 3.
@@ A F1 A F1 A F1 A F1 A F1 A F1 A F1
The results indicate no clear pattern regarding which rubric points can consistently be graded more accurately than others. Currently, Llemma and Llemma-based grading models perform well enough for use by students as practice for writing mathematical proofs. In future work, we want to investigate how to provide more accurate and detailed feedback to students, as the correctness of the models is not guaranteed.
@@
We randomly sampled 20 model-graded submissions and 15 human-graded submissions to P1 and P2 to qualitatively analyze the error patterns of the grading models and human graders. One key observation for grading models was the reduced accuracy when handling mathematical formulae. Some calculation mistakes went unnoticed by the grading model (for example, writing n=1 as the base case when it should be n=0), and some correct values were marked as incorrect (for example, correct base cases of n=0 and n=1 were identified as incorrect by the grader). On the other hand, human graders were able to grade the formulae correctly, but they sometimes ignore subtle details in the proof. For example, a correctly but implicitly written bound for the induction hypothesis was marked as incorrect by one of the recruited human graders. In addition, TA graders seem to have a hard time telling the difference between rubric points R3 (stating the inductive hypothesis) and R4 (setting the bound of the inductive hypothesis), marking them both wrong when only one of them was actually wrong. The grading model was more effective at making the distinction correctly
@@ 4.3 How large of a dataset do we need?
For all four models, the performance is still increasing even with 1100 training data points, so there is space for improvement by collecting more data. Model performance might also be better by making the grading models more complex, such as adding extra layers. However, most of the improvement is achieved with the 400 training size threshold. For Llemma34b, the performance at 400 training data already achieves 98.8% of the full data size accuracy. Based on these results, we conclude that training a model with 400 proofs is sufficient for adequate results.
@@ 5 User Study Methods
Using the trained grading pipelines, we implemented an autograder for proof by induction problems using the online assessment system PrairieLearn [51]. Due to computational resource restrictions and the comparable performance of GPT-3 to Llemma on the 3 proofs used in our user study, we chose to use GPT-3 as the embedding model for our user study. For each student-submitted proof on a problem, the grading pipeline evaluates the proof using the rubrics R1-R7 introduced in 3.1, producing 7 correct/incorrect results. The final score for each submission is the sum of the 7 rubric points, ranging between 0 and 7, and then rescaled to a percentage between 0-100. For each submission, we first show a general feedback based on the proof’s percentage score, such as “you are almost there” or “your proof needs more work”, encouraging students to revise their submissions. Additionally, each rubric point is mapped to a specific feedback sentence that is shown to the student when a mistake is identified for that rubric point. For instance, an error in rubric item R4 (setting the bound of the inductive hypothesis) generates the feedback: “It appears your inductive hypothesis is missing or incorrect.” According to the taxonomy of feedback introduced in Shute [6], the feedback provided by our grading pipeline is partially elaborated. We tell the students a rubric point which their proof is failing, but we do not explain to them how to fix it. Figure 4 shows an example page from PrairieLearn where students can submit their proofs and receive automatic feedback. We did not use generative models for feedback due to concerns about their accuracy at the time of the study design, and because we felt that rubric-point based feedback would be sufficiently elaborated for students to improve their proofs. As we will discuss more in Section 7.2, some students felt that the feedback was sufficiently helpful while others felt that it should have been more elaborated.
@@ 5.1 Study Setup
The learning activity consisted of three parts. In the first part, students were expected to read a 6-page textbook section about induction, including information on the theoretical ground for induction, the steps of an induction proof, and an example problem with solution. The second part asked students to complete 3 induction problems (P1-P3 as defined in 3.1). We only consider 3 problems out of the 4 used in the model training process for two reasons: first, we wanted to limit the length of the learning activity to 1 hour, which would not be reasonable if we expected students to read the text and complete all 4 problems; second, we removed P4 due to its low accuracy compared to P1-P3, as this could potentially hinder the students’ learning experiences. The third part of the learning activity included a feedback survey, with the following questions:
@@
The first subsection of the survey includes questions S01-S03, which asks the students about their prior experience with human graders in the course. The second subsection includes questions S04-S07, focusing on students’ perceptions of using AI for autograding. The last questions ask students to share their experiences with the AI autograder during the learning activity. The survey questions are presented on a five-point Likert scale using the following options and corresponding numeric values: “Strongly disagree” (-2), “Disagree” (1), “Neutral” (0), “Agree” (1), “Strongly agree” (2). Negatively worded questions (e.g., “I would prefer to wait for a week to get human grading feedback for the proofs I write for my homework than use the autograder for instant feedback”) are reverse coded in the statistical analysis.
To get a better understanding of student perceptions of the graders, we included the following open-ended survey questions:
@@ 5.2 Experimental Conditions
Both the First and Random groups had access to the AI autograders, forming our treatment groups. Students in the First group received feedback on the first rubric point (out of the seven) that the autograder identified. The order of these rubric points follows the logical sequence of R1–R7, corresponding to the typical flow of an induction proof. For example, an induction proof commonly begins with stating and proving the base cases, followed by stating the goal for the induction steps, and then completing the steps themselves. The strategy of reporting the first incorrect rubric point is based on some students’ preference to work incrementally, fixing the earlier parts of their work before moving on to the next part. When we designed the experiment, we expected this strategy to be the most helpful to students.
Students in the Random group received feedback for one of rubric points, randomly selected from the rubric items the autograder identified as incorrect. This approach was included in our study to explore whether different feedback strategies yield varying effects. The random strategy aimed to expose students to a broader range of feedback on the question, potentially helping them identify actionable steps to correct their submission. This approach could be especially beneficial for students who might otherwise remain stuck on the same part of the proof for an extended period of time.
@@ 6 User Study Results
6.1 RQ1: Are Students Able to Improve Their Proofs Using The Autograder?
@@ 6.2 Survey Results
Figure 8: Responses from survey questions given to only First and Random. The percentages shown on each distribution represent the percentage of negative responses (disagree & strongly disagree), the percentage of neutral responses, and the percentage of positive responses (agree & strongly agree).
The survey results are shown in Figure 7 and Figure 8. The three subsections of the surveys address different aspect of student perceptions: S01-S03 address student perceptions of human graders, S04-S07 address students willingness to have AI graders integrated into various parts of their course, and S08-S10 address student perceptions of the AI graders. We calculated a Cronbach’s $\alpha$ for each of the three subsections. The calculated Cronbach’s $\alpha$ are 0.81, 0.72, 0.82 respectively, indicating good internal reliability. The Self-eval group were not asked questions S08-S14 as these students did not interact with the autograder during the learning activity.
@@ 6.2.1 RQ2: Are students satisfied with the autograder and feedback?
We use responses from survey questions S01-S03 and S08-S10, as shown in Table  7, to address RQ2. S01-S03 capture students’ perceptions of human graders, while S08-S10 assess their perceptions of the AI autograder. Specifically, S01 and S08 measure the perceived accuracy of the graders, S02 and S09 evaluate their helpfulness, and S03 and S10 measure overall satisfaction with the graders. On average, human graders are perceived to be more accurate and helpful than the AI autograder, with statistically significant differences ($p<0.05$). However, although the mean satisfaction rating is higher for human graders, the differences are not statistically significant for either the First or Random groups. We hypothesize that this is due to the known benefits of the AI autograder, such as timely feedback and the ability to submit an unlimited number of times.
@@ 6.2.2 RQ3: Does using the autograder make students more willing to use AI autograders?
We aim to understand students’ perceptions of and trust in using an autograder for their coursework across various settings. Specifically, we want to determine whether interactions with the autograder can foster more positive perceptions. These are reflected through S04-S07 (see Table 8). S04 examines students’ preferences between receiving instant feedback from AI autograders and obtaining more reliable feedback from human graders. S05 assesses the level of trust students have in AI autograders in general. S06 and S07 explore potential applications of AI autograders.
@@ 7.2 Feedback Helpfulness
The elaborated nature of the feedback from our grader enabled students using the autograder to make good progress on the proofs, and being able to see the feedback almost instantly also provided students a better way of practicing writing the proofs.
@@
On the other hand, more elaborated feedback or instructions such as hints to fix the proofs would help student make even more improvements.
  • “Some of the feedback was basic. didn’t say anything specific. ”
Providing such automated and elaborated feedback should be possible with the help of generative models. We did not use generative models for this study due to concerns about low grading accuracy resulting in low feedback accuracy. At the time of the study, generative AI models stuggled to give feedback on long technical passages such as the induction proofs. However, they have improved over time so we can try them in future work. By leveraging few-shot training techniques such as Retrieval-Augmented Generation [52] and including the more accurate grading outputs from our models, we can fine-tune chatbots such as ChatGPT to provide suggestions or hints with reduced likelihood of hallucination. These fine-tuned chatbots can also better align with the specific learning objectives of individual courses. This potential enhancement to our system can be explored in future work, allowing us to get a better understanding into the quality of AI-generated feedback.
@@ 7.3 Perceptions on AI grading
Proof grading is a subjective task, which can lead to low levels of trust from students. When designing the learning activities, we initially hoped that the interactions with our AI autograder would increase students’ trust and willingness to use such tools in the future. However, results in 6.2 do not confirm our initial hypothesis. Students in First and Random are slightly more willing to use an AI autograder in various settings, but the difference is not statistically significant. Simply using the autograder does not sufficiently increase students’ willingness to use AI autograders in general. One potential strategy to build trust in AI autograders is to provide more detailed feedback and explanations during the grading process. Research by Ha and Kim [53] indicates that offering additional information about AI models can be effective in reducing cognitive biases and fostering trust. Incorporating such measures into future iterations of the autograder could improve its acceptance and perceived reliability among students.
@@ 8 Conclusions and Future Work
In this study, we developed a set of machine learning models to autograde freeform mathematical proofs. These models enable fast training without the need for extensive fine-tuning, and they achieve accuracy comparable to that of most human graders recruited for this study. These models can be used to develop NLP graders which could be deployed on large-scale educational platforms in the future, enabling more students to receive the necessary support when learning to write mathematical proofs.
After creating and deploying autograded proof by induction problems using the grading models, we conducted a user study to evaluate the autograder’s impact. Our results show that students who used the autograder for feedback to make revisions to their submissions scored over 10% higher on the proof problems than students who did not use the autograder. This is an encouraging sign for creating more autograded problems in the future. Despite being helped by the autograder, many students still preferred feedback from human graders, even if it was delayed.
In the future, fostering greater student trust in AI autograders will be an important area of focus. Efforts can be put on improving the grading models as well. Since our current models were trained exclusively on mathematical induction problems, their ability to grade other types of proof such as proof of bijection or geometry is still unknown. We plan to collect data on various types of mathematical proof problems, and check if our grading models can perform equally well on these problems.
Additionally, he cost of training and running these models can also be reduced. In our current design, GPT-3-based models seem to be the best choice as they can be scalable at a low cost and achieve a relatively good performance. However, there is still a lack of open-source language models that are computationally efficient, cost-effective, and capable of achieving comparable performance. Developing such models could allow for greater customization and broader application of AI autograders, making this an important focus for future work with more advanced language models.